OpenStudy (moongazer):

teach me on how to get the value of this 3 variables I have slight knowledge about this but i can't get the value with these systems of linear equations help me please Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (moongazer):

2x+3y-z=4 3x-4y+4z=11 5x-2y+z=5

OpenStudy (moongazer):

this is what i made: 2x+3y-z=4 (1) 3x-4y+4z=11 (2) 5x-2y+z=5 (3) I multiplied eq. 1 by 4 (2x+3y-z=4)4 then added it to eq.2 8x+12y-4z=16 (1) 3x-4y+4z=11 (2) =11x+8y=27 (4) then i multiplied eq.3 to -4 and added to eq.2 (5x-2y +z=5)4 then added it to eq.2 3x-4y+4z=11 eq.2 -20x+8y-4z=-20 eq. 3 =-17+4y=9 (5) then i multiplied (5) to -2 34x-8y=-18 (5) then i added it to (4) 11x+8y=27 34x-8y=-18 = 45x=9 x=9/45 i know my answer is wrong, but where did i go wrong?

OpenStudy (anonymous):

why do you say your answer is wrong?

OpenStudy (moongazer):

because i got zero in my quiz?

OpenStudy (moongazer):

here is the answer but i can't understand it http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e2109660b8ba00f8f002261

OpenStudy (anonymous):

8x+12y-4z=16 3x-4y+4z=11 --------------- + 11x + 8 y = 27 3x-4y+4z = 11 20x-8y+4z=20 --------------- - -17x+4y=-9 11x+8y = 27 -34x+8y=-18 -------------- - 45x = 45 x = 1 Your approach is correct, but I think you may have made a mistake with the + or - sign

OpenStudy (anonymous):

Check your arithmetic for equation (5)

OpenStudy (callisto):

then i multiplied eq.3 to -4 and added to eq.2 (5x-2y +z=5)4 then added it to eq.2 3x-4y+4z=11 eq.2 -20x+8y-4z=-20 eq. 3 =-17+4y=9 (5) it should be =-17+4y=-9 (5) I've replied this in http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e2109660b8ba00f8f002261